One (1) gram of aradioactive substance disintegrates at the rate of 3.7 × 1010 per

second. The atomic weight of the substance is 226. Calculate its mean life.


1
Expert's answer
2022-01-20T10:09:03-0500

"\tau=\frac NA=\frac{mN_A}{MA}=7.2\cdot 10^{10}~s."